### v2

# Nonlinear Drift Waves Beyond the Hasegawa-Mima Model

Abstract
A general nonlinear equation for drift waves is derived, incorporating both the nonlinear electron density term (scalar nonlinearity) and the ion vorticity term (vector nonlinearity). It is noted that the Hasegawa-Mima (HM) equation [A. Hasegawa and K. Mima, The Physics of Fluids 21, 87 (1978)] includes only the ion vorticity term, while it omits the nonlinear electron density term, which is a dominant nonlinear effect. The general nonlinear equation reduces to the Korteweg–de Vries (KdV) equation [H. Saleem, Physics of Plasmas 31, 112102 (2024)] if the normalized electrostatic potential is assumed to depend only on a single spatial coordinate along the direction of predominant wave propagation, i.e., $\Phi=\Phi(y)$. If the scalar nonlinearity is artificially neglected and a two-dimensional dependence of the electrostatic potential is assumed, $\Phi=\Phi(x,y)$, the equation reduces to a Hasegawa-Mima-like equation, which admits solution in the form of dipolar vortices. It is further pointed out that, in the presence of a Boltzmann density distribution for electrons, the HM equation was derived using the approximation $\frac{\partial_{t}\tilde{n}_{i}}{n_{i0}}\simeq\partial_{t}\Phi$, whereas the more accurate expression $\frac{\partial_{t}\tilde{n}_{i}}{n_{i0}}\simeq\partial_{t}(\Phi+\frac{1}{2}\Phi^{2})$ should be used.
preprint: AIP/123-QED
Key Words: Nonlinear Drift Waves, Korteweg de-Vries Equation, Dip Solitons, Hasegawa Mima Equation, Reductive Perturbation Method.
I Introduction
The drift wave is an important low-frequency electrostatic normal mode in magnetized inhomogeneous plasmas, due to its relevance to cross-field particle (and energy) transport and turbulence [1, 2, 3]. Since most magnetized space [4], astrophysical [5], and laboratory plasmas [2, 6] are inhomogeneous, the drift wave has attracted significant attention from researchers since its discovery [7, 8]. It is a stable mode in a collisionless cold-ion plasma under fluid theory, but becomes unstable when electron-ion collisions are considered. In collisional plasmas, it is known as the drift dissipative instability [9].
This wave is also referred to as the universal instability, as kinetic theory predicts that it can extract energy from electrons and become unstable through wave–particle interactions. A field-aligned electron current can further enhance this instability [10, 11], even in the absence of collisions. On the other hand, two-fluid theory shows that a field-aligned ion shear flow can give rise to unstable drift waves even in a cold ion collisionless plasma [12]. Drift waves have also been studied in hot plasmas using multi-fluid models to explore their role in the self-heating of the solar corona [13].
Nonlinear drift waves can give rise to vortices predicted by the Hasegawa-Mima (HM) equation [14], shocks [15], and solitons [16, 17]. Both linear and nonlinear drift waves have recently been investigated in multi-component space plasmas [18]. Several authors have attempted to explain solar coronal heating using drift waves in a kinetic framework applied to the hot solar corona [19, 20, 21].
Recently, it has been pointed out in [17] that the Korteweg–de Vries (KdV) equation for drift waves should be derived by normalizing the ion velocity with the electron diamagnetic drift speed, which is the intrinsic phase speed of the drift wave. For this purpose, the parallel component of ion velocity has been neglected, and the KdV equation has been derived using the reductive perturbation method (RPM). If parallel ion velocity is retained, the ion acoustic wave couples with the drift wave [22], and the resulting phase speed corresponds to the ion acoustic speed in the coupled linear dispersion relation. In earlier derivations of the KdV equation, the ion parallel motion was included, and ion velocity was normalized by the ion acoustic speed [16, 23]. The KdV–Burgers equation was also derived for collisional plasmas using this approach [24].
A key point emphasized in Ref. [17] is that the drift wave predominantly propagates in a direction perpendicular to both the density gradient and the ambient magnetic field, with a phase speed equal to the electron diamagnetic drift velocity. Therefore, the ion parallel motion should be neglected in order to accurately describe the behavior of nonlinear drift waves in the small-amplitude limit. Inertia-less electrons follow the magnetic field lines, and the scalar nonlinear density term is not canceled by the ion parallel current in the case of drift waves. This leads to the formation of density cavities [17], as opposed to the density humps predicted by the KdV equation for ion acoustic waves (IAWs) [25].
It is also important to highlight that the well-known Hasegawa-Mima (HM) equation [14] does not account for electron density nonlinearity arising from the Boltzmann distribution. Surprisingly, in the original formulation by Hasegawa and Mima, the nonlinear contribution from inertia-less Boltzmann electrons was neglected. To emphasize this point, we derive a general nonlinear equation for drift waves. This approach allows for a clear comparison of the magnitudes of the scalar and vector nonlinear terms. We demonstrate that the scalar nonlinearity, which emerges from the Boltzmann electron density distribution, generally dominates over the vector nonlinearity associated with ion vorticity. Furthermore, if the weaker nonlinear term is neglected under appropriate conditions, we recover the result presented in Ref [17]. Finally, we revisit the derivation of the HM equation to obtain a corrected nonlinear equation for drift waves.
II Nonlinear equation for drift waves
First, we just mention the simplest drift wave geometry in Cartesian coordinates to lay down the background for highlighting the missing scalar nonlinearity in Hasegawa-Mima (HM) equation [14]. If we choose the ambient magnetic field to be along z-axis ${\bf B}_{0}=B_{0}\hat{z}$, the density gradient along negative x-axis i.e. $\nabla n_{0}=-\hat{x}\mid\frac{dn_{0}}{dx}\mid$ and density gradient scale length $L_{n}=\kappa_{n}^{-1}=$ constant with $\kappa_{n}=\mid\frac{1}{n_{0}}\frac{dn_{0}}{dx}\mid$, then the linear drift wave propagates predominantly along y-direction with $k_{z}<<k_{y}$. Wave propagation is considered to be in yz-plane with wave vector ${\bf k}=(0,k_{y},k_{z})$. The small but non-zero component of the wave vector along field lines allows the inertia-less electrons to follow the Boltzmann density distribution in wave’s oscillating electrostatic potential. The conditions for this wave to propagate in a collision-less plasma can be summarized as follows,
$\omega\ll\Omega_{i};v_{Ti}k_{z}<<\omega<<v_{Te}k_{z};k_{z}\ll k_{y};\kappa_{n}<<k_{y}$ (1)
where $v_{Ti}^{2}=\frac{T_{i}}{m_{i}}$, $v_{Te}^{2}=\frac{T_{e}}{m_{e}}$, and $\Omega_{i}=\frac{eB_{0}}{m_{i}c}$. For Fourier analysis, we need the local approximation to be valid i.e. $\kappa_{n}\ll k_{y}$. Since its a low frequency wave, therefore the inertial-less electrons ($m_{e}\rightarrow 0$) are assumed to follow the Boltzmann distribution,
$\frac{n_{e}}{n_{e0}}\simeq\exp({\frac{e\varphi}{T_{e}}})\simeq 1+\varphi+\frac{1}{2}\varphi^{2}+...$ (2)
The perpendicular component of cold ion velocity in drift approximation $\omega<<\Omega$ is expressed in the following form,
${\bf v}_{i\perp}=-\frac{c}{B_{0}}\nabla_{\perp}\varphi\times\hat{z}-\frac{c}{B_{0}\Omega_{i}}(\partial_{t}+{\bf v}_{i}\cdot\nabla)\nabla_{\perp}\varphi={\bf v}_{E}+{\bf v}_{P}$ (3)
where ${\bf v}_{E}$ is the electric drift, ${\bf v}_{P}$ is ions polarization drift and $\nabla_{\perp}=(\partial_{x},\partial_{y})$. It gives,
$\nabla_{\perp}\cdot{\bf v}_{i\perp}\simeq-\frac{c}{B_{0}\Omega_{i}}(\partial_{t}+{\bf v}_{E}\cdot\nabla)\nabla_{\perp}^{2}\varphi$ (4)
where $\nabla_{\perp}\cdot{\bf v}_{E}=0$. Recently [17], it has been pointed out that the scalar nonlinearity which is represented by the term $\frac{1}{2}\varphi^{2}$ on right hand side (rhs) of relation (2) plays an important role in creating dip solitons. Using RPM, the KdV equation for drift waves has been obtained in the following form,
$\partial_{\tau}\Phi-\Phi\partial_{\xi}\Phi+\partial_{\xi}^{3}\Phi=0$ (5)
where $\Phi={\frac{e\varphi}{T_{e}}}$. The above KdV equation for drift waves has the form similar to the KdV equation for ion acoustic waves (IAWs) with an important difference that the nonlinear term in drift wave equation is negative and it can produce density cavities which have been observed in upper ionosphere [26, 27].
The KdV equation for drift waves has been obtained assuming that the physical quantities associated with this wave are functions of only one spatial coordinate i.e. $\varphi=\varphi(y)$ and under this situation, the Eq. (4) reduces to,
$\nabla_{\perp}\cdot{\bf v}_{i\perp}\simeq-\frac{c}{B_{0}\Omega_{i}}(\partial_{t})\nabla_{\perp}^{2}\varphi$ (6)
The above relation gives the dispersive term in KdV equation (5) and nonlinear term is the contribution of nonlinear electron density expressed in terms of electrostatic potential in Eq. (2).
Now we discuss the well-known Hasegawa-Mima (HM) equation for drift waves which admits the solutions in the form of dipolar vortex. Let us consider the ions continuity equation,
$\partial_{t}n_{i}+\nabla_{\perp}\cdot(n_{i}{\bf v}_{i})=0$ (7)
which can be expressed more explicitly as,
$(\partial_{t}+{\bf v}_{E}\cdot\nabla_{\perp})\frac{\tilde{n_{i}}}{n_{i0}}+\frac{\nabla n_{i0}}{n_{i0}}\cdot{\bf v}_{E}+\nabla_{\perp}\cdot{\bf v}_{P}$ (8)
$+[\frac{\tilde{n_{i}}}{n_{i0}}\nabla_{\perp}\cdot{\bf v}_{P}+\frac{\nabla n_{i0}}{n_{i0}}\cdot{\bf v}_{P}+\frac{\nabla\tilde{n_{i}}}{n_{i0}}\cdot{\bf v}_{P}]=0$
where $n_{i}=n_{i0}+\tilde{n_{i}}$ and $\tilde{n_{i}}$ is the perturbed part of ion density. If physical variables are assumed to be the functions of $(x,y)$ coordinates i.e. $\varphi=\varphi(x,y)$, then vorticity associated with drift wave dynamics becomes nonzero because in this case ×$\nabla\times{\bf v}_{E}=\frac{c}{B_{0}}\nabla_{\perp}^{2}\varphi(x,y)\hat{z}\neq 0$. The electric drift is larger than the polarization drift due to $\omega<<\Omega_{i}$. The terms in square bracket of Eq. (8) are smaller compared to the term $\nabla_{\perp}\cdot{\bf v}_{P}$ and hence it reduces to,
$(\partial_{t}+{\bf v}_{E}\cdot\nabla_{\perp})\frac{\tilde{n_{i}}}{n_{i0}}+\frac{\nabla n_{i0}}{n_{i0}}\cdot{\bf v}_{E}+\nabla_{\perp}\cdot{\bf v}_{P}\simeq 0$ (9)
Using Eqs. (3) and (9) along with Boltzmann distribution for electrons, the Hasegawa and Mima equation can be written in the following form,
$\partial_{t}(\Phi-\rho_{s}^{2}\nabla_{\perp}^{2}\Phi)+D_{e}(\nabla_{\perp}\Phi\times\hat{z}\cdot\nabla)[\rho_{s}^{2}\nabla_{\perp}^{2}\Phi+\ln\frac{1}{n_{i0}}]=0$ (10)
where $\rho_{s}=\frac{c_{s}}{\Omega_{i}}$.
Equation (10) is the same as Eq. (2.2.9) of Ref. [4]. This equation has been obtained by replacing $\frac{\hat{n_{i}}}{n_{i0}}$ with $\Phi$ in the first term on left hand side (lhs) of Eq. (9). Note that,
$(\partial_{t}+{\bf v}_{E}\cdot\nabla_{\perp})\frac{\tilde{n_{i}}}{n_{i0}}=(\partial_{t}+{\bf v}_{E}\cdot\nabla_{\perp})\Phi$ (11)
was considered and therefore the scalar nonlinear term of relation (2) i.e. $(\frac{1}{2}\varphi^{2})$ did not appear in the HM equation.
We try to obtain a general nonlinear equation for drift waves by employing RPM and this equation will reduce to KdV equation (5) in one dimensional case for $\varphi=\varphi(y)$. In two dimensional case $\Phi=\Phi(x,y)$, it will reduce to HM like equation if scalar nonlinear term is dropped forcefully.
To normalize Eq. (9), we define $n_{i}^{\prime}=\frac{n_{i}}{n_{0}}$, ${\bf v}_{E}^{\prime}=\frac{{\bf v}_{E}^{\prime}}{v_{De}}$, $y^{\prime}=\frac{y}{\rho_{s}}$, $x^{\prime}=\frac{x}{\rho_{s}}$, $\nabla^{\prime}=\rho_{s}\nabla$ and $t^{\prime}=\frac{c_{s}}{L_{n}}t$ where $v_{De}=D_{e}\kappa_{n}$ is the electrons diamagnetic drift velocity while $D_{e}=\frac{cT_{e}}{eB_{0}}$. Equation (9) in normalized form becomes,
$\partial_{t}(\frac{\tilde{n_{i}}}{n_{i0}}-\nabla_{\perp}^{2}\Phi)+(\nabla_{\perp}\Phi\times\hat{z}.\hat{x})+\frac{L_{n}}{\rho_{s}}(\nabla_{\perp}\Phi\times\hat{z}\cdot\nabla_{\perp})\nabla_{\perp}^{2}\Phi=0$ (12)
[…]
$\partial_{t}(e^{\Phi}-\nabla_{\perp}^{2}\Phi)+(\nabla_{\perp}\Phi\times\hat{z}.\hat{x})+\frac{L_{n}}{\rho_{s}}(\nabla_{\perp}\Phi\times\hat{z}\cdot\nabla_{\perp})\nabla_{\perp}^{2}\Phi=0$ (13)
We define the stretched coordinates $(\xi,\eta,\tau)$ following Washimi and Tanuiti [25] for $\epsilon<<1$,
$\xi=\epsilon^{1/2}(y-Ut);\eta=\epsilon^{1/2}x$ (14)
and
$\tau=\epsilon^{3/2}t$ (15)
which give $\partial_{y}=\epsilon^{1/2}\partial_{\xi}$, $\partial_{x}=\epsilon^{1/2}\partial_{\eta}$ and $\partial_{t}=(-\epsilon^{1/2}U\partial_{\xi}+\epsilon^{3/2}\partial_{\tau})$ where the wave phase speed $U_{0}$ is normalized with $v_{De}$ and hence $U=\frac{U_{0}}{v_{De}}=1$. Note that the ion velocity in normalized form can be expressed as,
${\bf v}_{i\perp}=-\frac{L_{n}}{\rho_{s}}\nabla_{\perp}\Phi\times\hat{z}-(\partial_{t}+{\bf v}_{E}\cdot\nabla_{\perp})\nabla_{\perp}\Phi={\bf v}_{E}+{\bf v}_{P}$ (16)
The expansion of physical quantities in terms of $\epsilon$ are defined as follows,
$n_{i}=1+\epsilon n_{i1}+\epsilon^{2}n_{i2}+...$ (17)
$\Phi=\epsilon\Phi_{1}+\epsilon^{2}\Phi_{2}+...$ (18)
$\frac{\tilde{n_{i}}}{n_{i0}}\simeq\epsilon\Phi_{1}+\epsilon^{2}(\frac{1}{2}\Phi_{1}^{2}+\Phi_{2})$ (19)
where quasi-neutrality $n_{i}=n_{e}$ has been used. Now we analyze each term of Eq. (12) in different order in stretched coordinates and write the expressions only up to order $\epsilon^{5/2}$ as follows.
$\partial_{t}\frac{\tilde{n_{i}}}{n_{i0}}\simeq(-\epsilon^{1/2}U\partial_{\xi}+\epsilon^{3/2}\partial_{\tau})[\epsilon\Phi_{1}+\epsilon^{2}(\frac{1}{2}\Phi_{1}^{2}+\Phi_{2})]$ (20)
$=\epsilon^{3/2}[-U\partial_{\xi}\Phi_{1}]+\epsilon^{5/2}[\partial_{\tau}\Phi_{1}-U\partial_{\xi}(\frac{1}{2}\Phi_{1}^{2}+\Phi_{2})]$
$\partial_{t}\nabla_{\perp}^{2}\Phi\simeq-(-\epsilon^{1/2}U\partial_{\xi}+\epsilon^{3/2}\partial_{\tau})(\epsilon\partial_{\eta}^{2}+\epsilon\partial_{\xi}^{2})(\epsilon\Phi_{1}+\epsilon^{2}\Phi_{2})$ (21)
$\simeq\epsilon^{5/2}[U(\partial_{\eta}^{2}+\partial_{\xi}^{2})\partial_{\xi}\Phi_{1}]$
$(\nabla_{\perp}\Phi\times\hat{z}.\hat{x})\partial_{y}\Phi=\epsilon^{1/2}\partial_{\xi}[\epsilon\Phi_{1}+\epsilon^{2}\Phi_{2}]=\epsilon^{3/2}[\partial_{\xi}\Phi_{1}]$ (22)
$+\epsilon^{5/2}[\partial_{\xi}\Phi_{2}]$
Let us look at the following term,
$\frac{L_{n}}{\rho_{s}}(\nabla_{\perp}\Phi\times\hat{z}\cdot\nabla_{\perp})\nabla_{\perp}^{2}\Phi=\frac{L_{n}}{\rho_{s}}(\partial_{y}\Phi\partial_{x}-\partial_{x}\Phi\partial_{y})\nabla_{\perp}^{2}\Phi$ (23)
$=\frac{L_{n}}{\rho_{s}}\{\Phi,\nabla_{\perp}^{2}\Phi\}$
where $\{\Phi,\nabla_{\perp}^{2}\Phi\}=(\partial_{y}\Phi\partial_{x}\nabla_{\perp}^{2}\Phi-\partial_{x}\Phi\partial_{y}\nabla_{\perp}^{2}\Phi)$ is Poisson bracket and,
$\{\Phi,\nabla_{\perp}^{2}\Phi\}=\epsilon^{1/2}\partial_{\xi}(\epsilon\Phi_{1}+\epsilon^{2}\Phi_{2})\epsilon^{1/2}\partial_{\eta}(\epsilon\partial_{\eta}^{2}+\epsilon\partial_{\xi}^{2})$ (24)
$(\epsilon\Phi_{1}+\epsilon^{2}\Phi_{2})-\epsilon^{1/2}\partial_{\eta}(\epsilon\Phi_{1}+\epsilon^{2}\Phi^{2})$
$\epsilon^{1/2}\partial_{\xi}(\epsilon\partial_{\eta}^{2}+\epsilon\partial_{\xi}^{2})(\epsilon\Phi_{1}+\epsilon^{2}\Phi_{2})$
In stretched coordinates, we define $\nabla_{\perp}=(\partial_{\eta},\partial_{\xi})$ and $\nabla_{\perp}^{2}=(\partial_{\eta}^{2}+\partial_{\xi}^{2})$ and hence the above relation in the lowest order becomes,
$\{\Phi,\nabla_{\perp}^{2}\Phi\}=\epsilon^{4}\{\Phi_{1},\nabla_{\perp}^{2}\Phi_{1}\}$ (25)
Therefore, in Eq. (12), we obtain,
$\frac{L_{n}}{\rho_{s}}(\nabla_{\perp}\Phi\times\hat{z}\cdot\nabla_{\perp})\nabla_{\perp}^{2}\Phi=\epsilon^{4}\frac{L_{n}}{\rho_{s}}\{\Phi_{1},\nabla_{\perp}^{2}\Phi_{1}\}$ (26)
Using the relations Eqs. (20, 21, 22) and ((26), Eq. (12) can be expressed in the following form,
$\epsilon^{3/2}[-U\partial_{\xi}\Phi_{1}+\partial_{\xi}\Phi_{1}]+\epsilon^{5/2}[\{\partial_{\tau}\Phi_{1}-U\partial_{\xi}(\frac{1}{2}\Phi_{1}^{2}+\Phi_{2})\}$ (27)
$+\partial_{\xi}\Phi_{2}+U\partial_{\xi}\nabla_{\perp}^{2}\Phi_{1}]+\epsilon^{4}\frac{L_{n}}{\rho_{s}}\{\Phi_{1},\nabla_{\perp}^{2}\Phi_{1}\}=0$
In the phase speed frame of the waves, if we assume $\frac{L_{n}}{\rho_{s}}$ of the order of $\epsilon^{-3/2}$, then in order $\epsilon^{5/2}$, we obtain,
[$[\{\partial_{\tau}\Phi_{1}-U\partial_{\xi}(\frac{1}{2}\Phi_{1}^{2}+\Phi_{2})\}+\partial_{\xi}\Phi_{2}+U\partial_{\xi}\nabla_{\perp}^{2}\Phi_{1}]$ (28)
$+\frac{L_{n}}{\rho_{s}}\{\Phi_{1},\nabla_{\perp}^{2}\Phi_{1}\}\simeq 0$
where $[-U\partial_{\xi}\Phi_{1}-\partial_{\xi}\Phi_{1}]=0$ for $U=1$ and the above equation reduces to,
$[\partial_{\tau}\Phi-\partial_{\xi}(\frac{1}{2}\Phi^{2})+\partial_{\xi}\nabla_{\perp}^{2}\Phi]+\frac{L_{n}}{\rho_{s}}\{\Phi,\nabla_{\perp}^{2}\Phi\}\simeq 0$ (29)
where $\Phi_{1}=\Phi$ has been used. The vorticity term can be of the same order as the scalar nonlinear term ($\frac{1}{2}\Phi^{2}$) if $\frac{L_{n}}{\rho_{s}}\simeq\epsilon^{-3/2}$ that is $\frac{\rho_{s}}{L_{n}}\simeq\epsilon^{3/2}$. Equation (29) is the general nonlinear equation for pure drift waves in stretched coordinates $(\eta,\xi,\tau)$. If $\epsilon^{4}\frac{L_{n}}{\rho_{s}}<<\epsilon^{5/2}$, then the Eq. (29) in order $\epsilon^{5/2}$ becomes,
$[\partial_{\tau}\Phi-\partial_{\xi}(\frac{1}{2}\Phi^{2})+\partial_{\xi}\nabla_{\perp}^{2}\Phi]\simeq 0$ (30)
The solution of nonlinear Eq. (29) is not straightforward because it contains the effects of compressibility via electron scalar nonlinear term [17] and ions vorticity term which was considered by Hasegawa and Mima [14]. The KdV equation can be obtained by simply ignoring the higher order vorticity term in Eq. (29). But Hasegawa-Mima like equation can be obtained only if the nonlinear electron density term is dropped forcefully. Both these limiting cases are presented in the following sections for the sake of completeness.
III Soliton solution
If $\Phi=\Phi(y,t)$ is assumed, then the vorticity term disappears in Eq. (29) and it reduces to Eq. (5) because $\nabla_{\perp}^{2}=\partial_{\xi}^{2}$ in this case. In moving coordinate $\chi=\xi-M\tau$ where $\frac{M_{0}}{v_{De}}$ is normalized speed of the solitary structure, the solution of Eq. (5) turns out to be,
$\Phi(\chi)=-\mid\Phi_{m}\mid Sech^{2}(\frac{\chi}{W})$ (31)
where the maximum normalized amplitude is $\mid\Phi_{m}\mid=3M<1$ and width of the solitary structure is $W=\sqrt{\frac{4}{M}}$. Equations (5) and (30) are the same as equations (34) and (36) of Ref. [17]. The above KdV equation creates dip soliton while the hump soliton is produced by the KdV equation for ion acoustic waves.
IV Vortex solution
If scalar nonlinearity is ignored, then Eq. (29) reduces to,
$\frac{\rho_{s}}{L_{n}}[\{\partial_{\tau}\Phi+\partial_{\xi}\nabla_{\perp}^{2}\Phi]-\{\nabla_{\perp}^{2}\Phi,\Phi\}=0$ (32)
In form, this is somewhat different from standard HM equation (10) because it has been obtained in stretched coordinates assuming $U=1$. However, it admits a solution in the form of dipolar vortex similar to HM equation. For simplicity we denote the planar coordinates $(\eta,\xi)$ by $(X,Y)$ in this section. Let us define a moving coordinate $Y=\xi-V\tau$ where $V=\frac{V_{0}}{v_{De}}$ and $V_{0}$ is the constant speed of the vortex which gives $\partial_{\xi}=\partial_{Y}$ and $\partial_{\tau}=-V\partial_{Y}$. Let $a_{0}=-\frac{\rho_{s}}{L_{n}}$ and $b_{0}=V\frac{\rho_{s}}{L_{n}}$. Then (32) can be expressed as,
$a_{0}\partial_{Y}\nabla_{\perp}^{2}\Phi+b_{0}\partial_{Y}\Phi+\{\nabla_{\perp}^{2}\Phi,\Phi\}=0$ (33)
The equation (32) can also be written as,
$\{(\nabla_{\perp}^{2}\Phi-\lambda^{2}\Phi),(\Phi+a_{0}X)\}=0$ (34)
where $\lambda^{2}=(-\frac{b_{0}}{a_{0}})>0$. The above equation is satisfied if $(\nabla_{\perp}^{2}\Phi-\lambda^{2}\Phi)$ is a function of $(\Phi+a_{0}X)$. Let us choose the simplest case,
$\nabla_{\perp}^{2}\Phi-\lambda^{2}\Phi=C_{0}(\Phi+a_{0}X)$ (35)
where $C_{0}$ is a constant to be determined. The above equation can be expressed as,
$\nabla_{\perp}^{2}\Phi-\mu^{2}\Phi=d_{0}X$ (36)
where $\mu^{2}=(C_{0}+\lambda^{2})$ and $d_{0}=C_{0}a_{0}$. Now we transform the above equation from $(X,Y)$ to polar coordinates $(r,\theta)$ with $X=r\cos\theta$, $Y=\sin\theta$, and $\theta=\tan^{-1}\frac{Y}{X}$. In polar coordinates, Eq. (36) is,
$\nabla_{\perp}^{2}\Phi-\mu^{2}\Phi=d_{0}(r\cos\theta)$ (37)
We consider a circle of radius $R_{0}$ and find the solutions of Eq. (37) in inner $r<R_{0}$ and outer $r>R_{0}$ regions of this circle following Ref. [28]. Let us use the variable separation technique by defining $\Phi(r,\theta)=\psi(r)f(\theta)$ and $f(\theta)=\cos\theta$. Equation (37) takes the form,
$r^{2}\partial_{r}^{2}\Phi+r\partial_{r}\Phi-(\mu^{2}r^{2}+1)\Phi=d_{0}r^{3}\cos\theta$ (38)
Let $s_{0}^{2}=-\mu^{2}$ and $r^{\prime}=s_{0}r$, then Eq. (37) becomes,
$r^{\prime 2}\frac{d^{2}}{dr^{\prime 2}}\psi(r^{\prime})+r^{\prime}\frac{d}{dr^{\prime}}\psi(r^{\prime})+(r^{\prime 2}-1)\psi(r^{\prime})=d_{0}\frac{r^{\prime 3}}{s_{0}^{3}}$ (39)
Solution of above equation in the inner region turns out to be,
$\Phi_{In}(r,\theta)=(P_{1}J_{1}(s_{0}r)+\frac{rd_{0}}{s_{0}^{2}})\cos\theta$ (40)
where $P_{1}$ is a constant and $J_{1}(s_{0}r)$ is Bessel function of order one. The outer solution of (37) turns out to be,
$\Phi_{out}=Q_{1}K_{1}(\lambda r)\cos\theta$ (41)
where $Q_{1}$ is a constant and $K_{1}(\lambda r)$ is modified Bessel function of order one. The continuity condition for $\Phi$ requires that the inner and outer solutions and their derivatives must be equal at $r=R_{0}$. Using continuity conditions for $\Phi$ at $r=R_{0}$, one obtains,
$P_{1}=-\frac{\lambda^{2}}{(\lambda^{2}+s_{0}^{2})}[\frac{R_{0}d_{0}}{s_{0}^{2}J_{1}(s_{0}R_{0})}]$ (42)
and,
$Q_{1}=\frac{s_{0}^{2}}{(\lambda^{2}+s_{0}^{2})}[\frac{R_{0}d_{0}}{s_{0}^{2}K_{1}(\lambda R_{0})}]$ (43)
The only unknown is now the parameter $s_{0}$ which can be estimated numerically using the relation,
$\frac{K_{1}(\lambda R_{0})}{K_{2}(\lambda R_{0})}=-\frac{s_{0}}{\lambda}\frac{J_{1}(s_{0}R_{0})}{J_{2}(s_{0}R_{0})}$ (44)
where $J_{2}(s_{0}R_{0})$ is the Bessel function of order two and $K_{2}(\lambda R_{0})$ is modified Bessel function of order two. The solution of HM equation has been discussed in several research articles and books.
V Summary
A nonlinear equation for two-dimensional drift waves has been derived to demonstrate that the Hasegawa-Mima (HM) equation does not fully capture the dynamics of these waves. Specifically, the electron nonlinear term [17] must appear alongside the ion vorticity term in the HM equation. The flaw in the original derivation by Hasegawa and Mima [14] is identified: the authors overlooked the role of the Boltzmann distribution of electrons by using the simplified relation $\frac{\tilde{n}_{i}}{n_{i0}}=\Phi$ to describe the nonlinear ion density response. However, the correct expression should be $\frac{{n}_{i}}{n_{i0}}=e^{\Phi}$ which yields $\frac{\tilde{n}_{i}}{n_{i0}}=(\Phi+\frac{1}{2}\phi^{2})$. This correction introduces a scalar nonlinear term, $\frac{1}{2}\Phi^{2}$, which must appear in Eq. (13) along with the vorticity term.
Interestingly, in the drift wave’s phase velocity frame, the scalar nonlinearity from electrons and the vector nonlinearity from ions (i.e., vorticity) are inherently coupled. Drift waves propagate predominantly in a direction perpendicular to both the background density gradient $\nabla n_{0}$ and the ambient magnetic field ${\bf B}_{0}$. If the electrostatic potential depends only on a single coordinate—as in the chosen geometry $\Phi=\Phi(y)$—then the electron nonlinearity leads to the formation of density cavities [17]. In contrast, if the electron nonlinear contribution is artificially suppressed and a two-dimensional potential $\Phi=\Phi(x,y)$ is assumed, the ion vorticity generates dipolar vortices, similar to the solutions of the Hasegawa-Mima equation.
The general nonlinear equation has been derived in the frame moving with the wave phase speed, with the normalized speed set to $U=1$. From the author’s perspective, the electron nonlinearity is the dominant effect and should not be neglected in the derivation of nonlinear drift wave equation.
The key conclusions of this work can be summarized as follows:
  • I.
    A general nonlinear equation for pure drift waves has been derived, incorporating both scalar and vector nonlinearities.
  • II.
    The nonlinear electron density term is generally dominant over the ion vorticity nonlinearity.
  • III.
    The nonlinear term $\frac{1}{2}\varphi^{2}$ is missing in the original Hasegawa-Mima equation [14] because they approximated $\frac{\partial_{t}\tilde{n_{i}}}{n_{i0}}=\partial_{t}\Phi$ whereas the quasi-neutrality condition yields $\frac{\partial_{t}\tilde{n_{i}}}{n_{i0}}\simeq\partial_{t}(\Phi+\frac{1}{2}\Phi^{2})$ assuming the inertia-less electrons to follow the Bltzmann distribution.
This work is expected to be valuable for future studies of nonlinear drift waves, which have wide-ranging applications in space, astrophysical, and laboratory plasmas.
VI Data Availability Statement
The data used for the preparation of the presented results is included in the text of this work.